OpenStudy (anonymous):

Explain how to check whether the answer is reasonable. Monica made 350 cookies for the bake sale. She sold them for $0.65 each. She sold of the cookies. How much money did Monica collect? A. Answer: $172.00. Explanation: Explanation: She sold of the cookies, so she sold 280 cookies. If she charged 50¢ per cookie, she would have collected $140.00. Because she sold the cookies for 65¢ each, she collected more than $140.00. B. Answer: $182.00. Explanation: If she sold all of the cookies,

OpenStudy (anonymous):

she sold 4/5 of cookies

OpenStudy (anonymous):

B. Answer: $182.00. Explanation: If she sold all of the cookies, she would have collected $227.50. Because she only sold 4/5 of the cookies, she collected less than $227.50. C. Answer: $207.50. Explanation: If she sold all of the cookies, she would have made $227.50. But she didn't sell 1/5 of the cookies. If 1/5 is subtracted from $227.50, the answer is about $207.00.

OpenStudy (anonymous):

A 4/5 sold thats missing

OpenStudy (anonymous):

C? i think

OpenStudy (phi):

so the dollars should be 350 * 0.8 * 0.65 or 350 * 4/5 * 65/100 the second way we can do in our head 350/5 = 70 4*70= 280 280*65/100 28 * 65/10 14 * 65/5 14*13 169+13= 182

OpenStudy (phi):

though a calculator is faster

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Mehek14